OpenStudy (anonymous):

y(x)=22.5(x+9)-81 linear equation y(x)=22.5(x+9)-81 linear equation @Mathematics

OpenStudy (mathmagician):

y(x)=22.5x+202.5-81=22.5x+121.5

OpenStudy (anonymous):

then how do i get my y ponts and x ponts?

OpenStudy (mathmagician):

you choose one x value and then calculate the value of y, e.g. when x=0, y=22.5*0+121.5=121.5

OpenStudy (anonymous):

divide 121.5/22.5 for y